# Causally Reliable Concept Bottleneck Models

### v3

@@ 5.4 Fairness
C2BMs permit interventions to meet fairness requirements (Tab. 3).
We measure the Causal Concept Effect (CaCE) (goyal2019) of Attractive on Should be Hired before and after blocking the only path between them, i.e., performing an intervention on Qualified. Tab. 3 shows that C2BM is the only model able to successfully remove the influence, achieving a post-intervention CaCE of 0.0%. This difference stems from the model architectures: in CBM, CEM, and SCBM, all concepts are directly connected to the task, meaning interventions on one concept cannot block the influence of others. In contrast, C2BM enforces a structured causal bottleneck allowing for interventions to fully override the effects of parent nodes and block information propagation through the intervened node. This highlights C2BM’s ability to enforce causal fairness by eliminating biased pathways.
@@ 6 Conclusions
Future directions include a deeper investigation of OOD generalization with C2BMs, an extensive exploration of their role in causal inference (e.g, counterfactual queries extending , see App. A), and the identification of optimal intervention policies. Moreover, incorporating PAGs (Zhang2008) would enable modeling of hidden confounders.
@@ A.1 Pearl’s framework of causality
Contemporary research in causal inference and causal machine learning predominantly builds on the framework introduced by Pearl2009 (see also Pearl2019). This framework centers on an agent’s ability to reason about underlying causal mechanisms, going beyond mere statistical associations observed in data. Pearl formalizes this capacity through the notion of answering different types of what-if questions, structured into a three-level hierarchy (see Figure 6).
Figure 6: Pearl’s Hierarchy as represented in Pearl2018.
@@
While standard machine learning models  [^5] This excludes models specifically designed for causal inference, such as those in the field of causal machine learning (kaddour2022causal). can generally handle only observational questions, reasoning at interventional and counterfactual levels necessitates dedicated causal models and inference methods (Pearl2009; peters2017elements).
@@ Appendix B Concept discovery details
CBMs-like architectures tipically rely on labeled data for each concept, a costly process that can hinder their practical adoption. Label-free Concept Bottleneck Models (Label-free CBMs) (oikarinen2023) address this issue by automatically generating concepts and assign concept labels using pre-trained models.
In our paper, we apply an approach similar to the one followed in oikarinen2023 to the Siim-Pneumothorax dataset, which lacks concepts and concept annotations. Using GPT-4o, we first generate candidate concepts with a specific prompt (see App. C.3), and apply a multi-stage filtering process:
@@
  • filter out concepts too similar to class labels or to each other. Specifically, we use the CXR-CLIP model – a CLIP-based model pretrained on medical imaging datasets (johnson2019; irvin2019; wang2017; you2023) – to encode both concepts and class labels, and discard any concept with cosine similarity > 0.9 to either a class label or another concept;
@@ C.1 Causal discovery algorithms
Algorithms for addressing the causal discovery problem can be broadly classified into two main categories (peters2017elements): [^6] Here, we restrict our discussion to causal discovery algorithms that assume the set of observed variables sufficiently captures the relevant causal influences. For more complex scenarios, see, e.g., (peters2017elements).
@@
  • Score-based methods. These methods define a scoring function over potential graph structures and search for the graph that maximizes it, often using criteria such as the Bayesian Information Criterion (BIC) (peters2017elements). The results from score-based methods are often comparable to those from independence-based approaches, as graphs that violate conditional independencies tend to result in poor model fits.
For our experiments, we adopted the Greedy Equivalence Search (GES) algorithm (chickering2002), a score-based method that performed well in our setting, as shown in Table 8 in App. G.3.
@@
In our implementation, we use the GES algorithm provided by the causal-learn Python library (zheng2024), employing the Bayesian Dirichlet equivalent uniform (BDeu) scoring criterion for discrete variables (heckerman1995learning; chickering2002).
@@ C.2 LLM and RAG
LLMs (brown2020language; jiang2024mixtral) are capable of answering complex queries without additional training. However, they can be unreliable and prone to hallucinations (ji2023survey; zhang2023siren). Retrieval Augmented Generation (RAG) (lewis2020retrieval) addresses this by retrieving relevant textual information and appending it to the query, thereby improving the accuracy and reliability of LLM outputs. In the context of structural learning, LLMs have been utilized to construct causal graphs by employing ad-hoc prompts specifically designed to condition the model for answering causal queries (antonucci2023zero; long2023can; zhang2024causal). To enhance the causal graph discovery process, we utilize an LLM integrated with a RAG to either direct or eliminate edges that remain undirected by the causal discovery algorithm. The information retrieval process is outlined as follows:
@@
    Ranking: The causal query is transformed by the LLM using a query transformation approach (gao2023retrieval), which improves the semantic alignment of the query with the relevant document chunks. After transformation, all the retrieved chunks and the modified causal query are processed by a sentence transformer (reimers2019sentence), specifically the multi-qa-mpnet-base-dot-v1 model. At this stage, cosine similarity is computed between the embedded transformed causal query and each embedded chunk.
@@ Appendix E Dataset details
The MNIST dataset (lecun2010) is a large collection of freely available grayscale images of handwritten digits. It consists of $60000$ training images and $10000$ test images, both drown from the same distribution. Each image is labeled with the digit it represents. For our experiments, we download the original train and test dataset using the torchvision library (marcel2010) and reserve $10\%$ of the training set for validation. Additionally, we colorize each image based on its digit. Specifically, in the in-distribution setting (experiment in Sec.5.1), color is randomly assigned to each image in the training, validation, and test sets with equal probability over red, green, and blue. In the out-of-distribution setting (experiment in Sec.5.3), images in the training and validation sets with odd digits are colored green, while the remaining images are colored red or blue with equal probability. In the test set, images with even digits are colored green, and the remaining ones are colored red or blue with equal probability. For both the versions of this dataset, the following concepts are considered: Number, Color, and Parity (task). Finally, the images are preprocessed using a pre-trained ResNet-18 model with default weights from the torchvision library.
@@ E.2 Bayesian networks
For our experiments, we use synthetic datasets sampled from discrete Bayesian networks available in the bnlearn repository (https://www.bnlearn.com/bnrepository/). A Bayesian network is a probabilistic graphical model consisting of a DAG, where nodes correspond to random variables, and each node is associated with a conditional probability distribution (CPD). This CPD defines the probability of the node’s value, given the values of its parent nodes in the network (sharma2020). From the bnlearn repository, we select Bayesian networks with different dimensions and domains: Asia (asia), a small network focused on lung disease with $8$ nodes and $8$ edges; Sachs (sachs2005causal), a widely-used network modeling the relationships between protein and phospholipid expression levels in human cells with $11$ nodes and $17$ edges; Insurance (binder1997), a network for evaluating car insurance risks with $27$ nodes and $52$ edges; Alarm (alarm), a network designed to provide an alarm message system for patient monitoring with $37$ nodes and $46$ edges; Hailfinder (abramson1996), a network designed to forecast severe summer hail in northeastern Colorado with $56$ nodes and $66$ edges. For each network, we generate $10000$ samples and create training, validation, and test datasets using a $70\%-10\%-20\%$ split.
@@ E.3 CelebA
CelebA (liu2015) is a large-scale face attributes dataset with more than $200.000$ celebrity images divided into training, validation and test set with $40$ binary attribute annotations. For our experiments, we first downloaded all the splits from the project website https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html.
@@ E.4 Siim-pneumothorax
This dataset is derived from the publicly available chest radiograph dataset provided by the National Institutes of Health (NIH) and it contains chest x-ray images with binary annotations indicating the presence or the absence of Pneumothorax. For our experiments, we use in particular the training annotations available on Kaggle (https://www.kaggle.com/competitions/siim-acr-pneumothorax-segmentation) and the corresponding training images from https://www.kaggle.com/datasets/abhishek/siim-png-images. The dataset is then split into training, validation, and test sets using a traditional $70\%-10\%-20\%$ partition. As it does not include concept annotations, we followed a procedure inspired by the methodology in oikarinen2023 to generate concepts and their corresponding annotations. More details can be found in App. B.
@@ E.5 CUBC
This dataset is derived from the publicly available Caltech-UCSD Birds-200-2011 (CUB) (he2019fine) dataset, which is widely used in the CBM community (koh2020; zarlenga2022concept). It contains 11,788 images across 200 bird categories, with 5,994 images for training and 5,794 images for testing. Each image is annotated in detail, including 312 binary attributes. For our experiments, we downloaded the dataset from https://data.caltech.edu/records/65de6-vp158 and further split the training set such that $10\%$ is used for validation. We then selected the 112 most frequently activated binary attributes to serve as our concepts of interest, as considered in zarlenga2022concept.
@@ Appendix F Experimental details
Python code and instructions for reproducing results across all datasets and methods are available within the code provided alongside the submission as supplementary material. We detail below the configurations and hyperparameters used for models’ instantiation and training. For a fair comparison across concept-based models, we standardize the concept encoder to a single-hidden-layer MLP. All models are trained using the Adam optimizer (kingma2014adam) for a maximum of 500 epochs, with early stopping based on a 30-epoch patience. We employ LeakyReLU as the activation function throughout.
The batch size is set to 512 for most datasets, with the exception of Siim-pneumothorax and SCBM, where it is reduced to 128 due to memory constraints. Following koh2020, we regularize the task loss to encourage concept learning, using a weighted sum of task and concept losses:
@@
Additionally, we apply random training-time interventions as proposed by zarlenga2022concept, with an intervention probability of 0.25. Regarding SCBM, we used the authors’ implementation at https://github.com/mvandenhi/SCBM. More precisely, we implemented the global variation using the configuration proposed by the authors.
@@ G.3 Ablation study on Causal Discovery
In this section, we evaluate the sensitivity of the causal discovery component in our pipeline to the choice of method for constructing a causal graph. To do so, we compare our method of choice for causal discovery, i.e., Greedy Equivalence Search (GES) (chickering2002), with different widely-used causal discovery algorithms that recover a CPDAG. Each of these methods is evaluated with and without refinement via the retrieval-augmented generation (RAG)-enhanced language model (LLM) employed in our study. Furthermore, we compare all these methods against the use of the LLM alone, as well as our retrieval-augmented LLM (LLM + RAG) applied directly to discover the whole graph. Specifically, the LLM is prompted with the causal prompt shown in App. C.3, with additional retrieved context appended to the query in the LLM+RAG setting.
@@
  • PC Algorithm (spirtes2000causation) (independence-based): A classical independence-based method that first estimates the undirected structure of the causal relations through a sequence of conditional independence tests, then orients edges using a set of predefined orientation rules (colombo2014). In our implementation, we use the PC algorithm provided by the causal-learn library (zheng2024), with a chi-squared independence test and a significance level of 0.05;
  • Fast GES (ramsey2017) (FGES) (score-based): A computationally efficient variant of GES that improves performance by storing intermediate evaluations and parallelizing expensive operations, enabling its application to large datasets (andrews2019; ramsey2017). In our implementation, we use the FGES algorithm provided by the py-tetrad library (ramsey2023py), employing the Bayesian Dirichlet Equivalent Uniform (BDeu) score for structure evaluation (heckerman1995learning).
@@ G.4 Ablation study on LLM type
  • Few-shot prompting: proposed in ye2024investigating, combines instruction prompting with a few examples;
  • Chain-of-Thought (CoT): proposed in wei2022chain, encourages the model to generate intermediate logical steps before generating the answer. This is the approach we used in the original manuscript.

### v4

@@ 5.4 Fairness
C2BMs permit interventions to meet fairness requirements (Tab. 3).
We measure the Causal Concept Effect (CaCE) (Goyal et al., 2019) of Attractive on Should be Hired before and after blocking the only path between them, i.e., performing an intervention on Qualified. Tab. 3 shows that C2BM is the only model able to successfully remove the influence, achieving a post-intervention CaCE of 0.0%. This difference stems from the model architectures:
@@
in CBM, CEM, and SCBM, all concepts are directly connected to the task, meaning interventions on one concept cannot block the influence of others. In contrast, C2BM enforces a structured causal bottleneck allowing for interventions to fully override the effects of parent nodes and block information propagation through the intervened node. This highlights C2BM’s ability to enforce causal fairness by eliminating biased pathways.
@@ 6 Conclusions
Future directions include a deeper investigation of OOD generalization with C2BMs, an extensive exploration of their role in causal inference (e.g, counterfactual queries extending , see App. A), and the identification of optimal intervention policies. Moreover, incorporating PAGs (Zhang, 2008) would enable modeling of hidden confounders.
@@ A.1 Pearl’s framework of causality
Contemporary research in causal inference and causal machine learning predominantly builds on the framework introduced by Pearl (2009) (see also Pearl (2019)). This framework centers on an agent’s ability to reason about underlying causal mechanisms, going beyond mere statistical associations observed in data. Pearl formalizes this capacity through the notion of answering different types of what-if questions, structured into a three-level hierarchy (see Figure 6).
Figure 6: Pearl’s Hierarchy as represented in Pearl and Mackenzie (2018).
@@
While standard machine learning models  [^5] This excludes models specifically designed for causal inference, such as those in the field of causal machine learning (Kaddour et al., 2022). can generally handle only observational questions, reasoning at interventional and counterfactual levels necessitates dedicated causal models and inference methods (Pearl, 2009; Peters et al., 2017).
@@ Appendix B Concept discovery details
CBMs-like architectures tipically rely on labeled data for each concept, a costly process that can hinder their practical adoption. Label-free Concept Bottleneck Models (Label-free CBMs) (Oikarinen et al., 2023) address this issue by automatically generating concepts and assign concept labels using pre-trained models.
In our paper, we apply an approach similar to the one followed in Oikarinen et al. (2023) to the Siim-Pneumothorax dataset, which lacks concepts and concept annotations. Using GPT-4o, we first generate candidate concepts with a specific prompt (see App. C.3), and apply a multi-stage filtering process:
@@
  • filter out concepts too similar to class labels or to each other. Specifically, we use the CXR-CLIP model – a CLIP-based model pretrained on medical imaging datasets (Johnson et al., 2019; Irvin et al., 2019; Wang et al., 2017; You et al., 2023) – to encode both concepts and class labels, and discard any concept with cosine similarity > 0.9 to either a class label or another concept;
@@ C.1 Causal discovery algorithms
Algorithms for addressing the causal discovery problem can be broadly classified into two main categories (Peters et al., 2017): [^6] Here, we restrict our discussion to causal discovery algorithms that assume the set of observed variables sufficiently captures the relevant causal influences. For more complex scenarios, see, e.g., (Peters et al., 2017).
@@
  • Score-based methods. These methods define a scoring function over potential graph structures and search for the graph that maximizes it, often using criteria such as the Bayesian Information Criterion (BIC) (Peters et al., 2017). The results from score-based methods are often comparable to those from independence-based approaches, as graphs that violate conditional independencies tend to result in poor model fits.
For our experiments, we adopted the Greedy Equivalence Search (GES) algorithm (Chickering, 2002), a score-based method that performed well in our setting, as shown in Table 8 in App. G.3.
@@
In our implementation, we use the GES algorithm provided by the causal-learn Python library (Zheng et al., 2024), employing the Bayesian Dirichlet equivalent uniform (BDeu) scoring criterion for discrete variables (Heckerman et al., 1995; Chickering, 2002).
@@ C.2 LLM and RAG
LLMs (Brown et al., 2020; Jiang et al., 2024) are capable of answering complex queries without additional training. However, they can be unreliable and prone to hallucinations (Ji et al., 2023; Zhang et al., 2023). Retrieval Augmented Generation (RAG) (Lewis et al., 2020) addresses this by retrieving relevant textual information and appending it to the query, thereby improving the accuracy and reliability of LLM outputs. In the context of structural learning, LLMs have been utilized to construct causal graphs by employing ad-hoc prompts specifically designed to condition the model for answering causal queries (Antonucci et al., 2023; Long et al., 2023; Zhang et al., 2024). To enhance the causal graph discovery process, we utilize an LLM integrated with a RAG to either direct or eliminate edges that remain undirected by the causal discovery algorithm. The information retrieval process is outlined as follows:
@@
    Ranking: The causal query is transformed by the LLM using a query transformation approach (Gao et al., 2023), which improves the semantic alignment of the query with the relevant document chunks. After transformation, all the retrieved chunks and the modified causal query are processed by a sentence transformer (Reimers, 2019), specifically the multi-qa-mpnet-base-dot-v1 model. At this stage, cosine similarity is computed between the embedded transformed causal query and each embedded chunk.
@@ Appendix E Dataset details
The MNIST dataset (LeCun et al., 2010) is a large collection of freely available grayscale images of handwritten digits. It consists of $60000$ training images and $10000$ test images, both drown from the same distribution. Each image is labeled with the digit it represents. For our experiments, we download the original train and test dataset using the torchvision library (Marcel and Rodriguez, 2010) and reserve $10\%$ of the training set for validation. Additionally, we colorize each image based on its digit. Specifically, in the in-distribution setting (experiment in Sec.5.1), color is randomly assigned to each image in the training, validation, and test sets with equal probability over red, green, and blue. In the out-of-distribution setting (experiment in Sec.5.3), images in the training and validation sets with odd digits are colored green, while the remaining images are colored red or blue with equal probability. In the test set, images with even digits are colored green, and the remaining ones are colored red or blue with equal probability. For both the versions of this dataset, the following concepts are considered: Number, Color, and Parity (task). Finally, the images are preprocessed using a pre-trained ResNet-18 model with default weights from the torchvision library.
@@ E.2 Bayesian networks
For our experiments, we use synthetic datasets sampled from discrete Bayesian networks available in the bnlearn repository (https://www.bnlearn.com/bnrepository/). A Bayesian network is a probabilistic graphical model consisting of a DAG, where nodes correspond to random variables, and each node is associated with a conditional probability distribution (CPD). This CPD defines the probability of the node’s value, given the values of its parent nodes in the network (Sharma et al., 2020). From the bnlearn repository, we select Bayesian networks with different dimensions and domains: Asia (Lauritzen and Spiegelhalter, 1988), a small network focused on lung disease with $8$ nodes and $8$ edges; Sachs (Sachs et al., 2005), a widely-used network modeling the relationships between protein and phospholipid expression levels in human cells with $11$ nodes and $17$ edges; Insurance (Binder et al., 1997), a network for evaluating car insurance risks with $27$ nodes and $52$ edges; Alarm (Beinlich et al., 1989), a network designed to provide an alarm message system for patient monitoring with $37$ nodes and $46$ edges; Hailfinder (Abramson et al., 1996), a network designed to forecast severe summer hail in northeastern Colorado with $56$ nodes and $66$ edges. For each network, we generate $10000$ samples and create training, validation, and test datasets using a $70\%-10\%-20\%$ split.
@@ E.3 CelebA
CelebA (Liu et al., 2015) is a large-scale face attributes dataset with more than $200.000$ celebrity images divided into training, validation and test set with $40$ binary attribute annotations. For our experiments, we first downloaded all the splits from the project website https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html.
@@ E.4 Siim-pneumothorax
This dataset is derived from the publicly available chest radiograph dataset provided by the National Institutes of Health (NIH) and it contains chest x-ray images with binary annotations indicating the presence or the absence of Pneumothorax. For our experiments, we use in particular the training annotations available on Kaggle (https://www.kaggle.com/competitions/siim-acr-pneumothorax-segmentation) and the corresponding training images from https://www.kaggle.com/datasets/abhishek/siim-png-images. The dataset is then split into training, validation, and test sets using a traditional $70\%-10\%-20\%$ partition. As it does not include concept annotations, we followed a procedure inspired by the methodology in Oikarinen et al. (2023) to generate concepts and their corresponding annotations. More details can be found in App. B.
@@ E.5 CUBC
This dataset is derived from the publicly available Caltech-UCSD Birds-200-2011 (CUB) (He and Peng, 2019) dataset, which is widely used in the CBM community (Koh et al., 2020; Zarlenga et al., 2022). It contains 11,788 images across 200 bird categories, with 5,994 images for training and 5,794 images for testing. Each image is annotated in detail, including 312 binary attributes. For our experiments, we downloaded the dataset from https://data.caltech.edu/records/65de6-vp158 and further split the training set such that $10\%$ is used for validation. We then selected the 112 most frequently activated binary attributes to serve as our concepts of interest, as considered in Zarlenga et al. (2022).
@@ Appendix F Experimental details
Python code and instructions for reproducing results across all datasets and methods are available within the code provided alongside the submission as supplementary material. We detail below the configurations and hyperparameters used for models’ instantiation and training. For a fair comparison across concept-based models, we standardize the concept encoder to a single-hidden-layer MLP. All models are trained using the Adam optimizer (Kingma and Ba, 2015) for a maximum of 500 epochs, with early stopping based on a 30-epoch patience. We employ LeakyReLU as the activation function throughout.
The batch size is set to 512 for most datasets, with the exception of Siim-pneumothorax and SCBM, where it is reduced to 128 due to memory constraints. Following Koh et al. (2020), we regularize the task loss to encourage concept learning, using a weighted sum of task and concept losses:
@@
Additionally, we apply random training-time interventions as proposed by Zarlenga et al. (2022), with an intervention probability of 0.25. Regarding SCBM, we used the authors’ implementation at https://github.com/mvandenhi/SCBM. More precisely, we implemented the global variation using the configuration proposed by the authors.
@@ G.3 Ablation study on Causal Discovery
In this section, we evaluate the sensitivity of the causal discovery component in our pipeline to the choice of method for constructing a causal graph. To do so, we compare our method of choice for causal discovery, i.e., Greedy Equivalence Search (GES) (Chickering, 2002), with different widely-used causal discovery algorithms that recover a CPDAG. Each of these methods is evaluated with and without refinement via the retrieval-augmented generation (RAG)-enhanced language model (LLM) employed in our study. Furthermore, we compare all these methods against the use of the LLM alone, as well as our retrieval-augmented LLM (LLM + RAG) applied directly to discover the whole graph. Specifically, the LLM is prompted with the causal prompt shown in App. C.3, with additional retrieved context appended to the query in the LLM+RAG setting.
@@
  • PC Algorithm (Spirtes et al., 2000) (independence-based): A classical independence-based method that first estimates the undirected structure of the causal relations through a sequence of conditional independence tests, then orients edges using a set of predefined orientation rules (Colombo et al., 2014). In our implementation, we use the PC algorithm provided by the causal-learn library (Zheng et al., 2024), with a chi-squared independence test and a significance level of 0.05;
  • Fast GES (Ramsey et al., 2017) (FGES) (score-based): A computationally efficient variant of GES that improves performance by storing intermediate evaluations and parallelizing expensive operations, enabling its application to large datasets (Andrews et al., 2019; Ramsey et al., 2017). In our implementation, we use the FGES algorithm provided by the py-tetrad library (Ramsey and Andrews, 2023), employing the Bayesian Dirichlet Equivalent Uniform (BDeu) score for structure evaluation (Heckerman et al., 1995).
@@ G.4 Ablation study on LLM type
  • Few-shot prompting: proposed in Ye et al. (2024), combines instruction prompting with a few examples;
  • Chain-of-Thought (CoT): proposed in Wei et al. (2022), encourages the model to generate intermediate logical steps before generating the answer. This is the approach we used in the original manuscript.